The pressure gauge is used the pressure of water flow at point A. If the pressure at B is 87 kPa, estimate the pressure at A in kPa. Assume all fluids are at 20 degrees celsius. ans: PA=96.4 kPa

have answer, need solution


Fig. 9 is used the pre t A in kPa. Assume all fluids are at 20°C. sed the pressure of water flow at point A. IH SAE 30 oil Gage B 6 cm Mercury 5 cm Water flow 1t cm 4 cm Fig. 9
Answer #1

Hello,

1. Density of SAE 30 oil at 20o degree C (do)= 878.7 kg/m3

Density of water (dw)= 1000 kg/m3

Density of water (dm)= 13600 kg/m3

SAE 30 oil Gage B 6 cm Mercury 7 cm 5 cm Water flow Datum Line 11 cm 4 cm Fig. 9

from the modiefied figure,

applying pressure balance in both limbs at datum level,

pressure in left limb= pressure in right limb

PA+ dw*g*hw = PB + do*g*ho + dm*g*hm

here, hw= 5 cm, ho= 6cm, hm= 7cm and taking g= 9.8 m/s2

PA = PB + do*g*ho + dm*g*hm - dw*g*hw

PA= 87000+ 878.7*9.8*0.06 + 13600*9.8*0.07- 1000*9.8*0.05

= 96.4 kPa

.................... Thank You